OpenStudy (anonymous):

Use the initial condition to find the particular solution of the differential equation. Differential Equation Initial Condition yy' − ex = 0 y = 8 when x = 0

OpenStudy (anonymous):

i don't really know this stuff, but can't you start with \[\int y y' =\int e^xdx\]?

OpenStudy (anonymous):

As satellite mentioned, you can separate the variables and integrate: \[\int y\,dy=\int e^x\,dx\] (Close enough :P)